OpenStudy (anonymous):

Find the derivative of the function h(x) = ln(9x+1/4x-2)

zepdrix (zepdrix):

\[\huge \log(\frac{ a }{ b })=\log(a)-\log(b)\] Use this rule of logarithms to split your problem into two separate logs. \[a=9x+1\]\[b=4x-2\]

OpenStudy (anonymous):

ok, then i would just have to subtract a from b to get an answer?

zepdrix (zepdrix):

No, we haven't taken a derivative. Applying the log rule just makes it EASIER to take a derivative. But you still need to find the derivative of the logs.

OpenStudy (anonymous):

i see. so a' = 9 while b' = 4, correct?

zepdrix (zepdrix):

Oh when you apply the chain rule? yes i suppose thats true :O don't take the derivative with a's and b's though XD I was just stating a rule for u hehe

zepdrix (zepdrix):

\[\large \ln(\frac{ 9x+1 }{ 4x-2 })=\ln(9x+1)-\ln(4x-2)\] I was just suggesting that you use that rule to write your problem like this.

OpenStudy (anonymous):

oh alright

OpenStudy (anonymous):

@zepdrix , so my next step would be finding the derivative of both equations?

zepdrix (zepdrix):

Both pieces? yes c: Remember how to take the derivative of the natural log, don't forget the chain rule! :D

OpenStudy (anonymous):

so the derivative for the fist piece is 9 and second is 4, what would i do with the chain rule

zepdrix (zepdrix):

 Understand how to do the first piece? :o